INAF queue observing, 2018 September 26-27

Observer: R. Carini, A. Gargiulo
Telescope Operator: D. G. Huerta
Support Astronomer: B. Rothberg on call
Instrument: LUCI

Summary

At the sunset the sky is clear
instruments: MODS2 red grating is not working
LBC-blue has some issues LUCI1 AO doesn't work.

Night duration (12\xB0 el twilight) 2:30- 12:15 UT = 9 h 45m

Observing time: 7h 5m (73%)

Weather time loss: h m (0%)

technical time loss 2h 40m (27%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Pinna Lo Peg LUCI2-AO FeII, K, H2 15 m / 15 m luci2.0155-0206 seeing between 1"-1.2"
telluric Antoniucci HIP116229 LUCI1+LUCI2 zj spec, Hkspec   luci1.0156-0163, luci2.0212-0219  
Antoniucci_11 V733 LUCI1+LUCI2 zj spec, Hkspec 0.4/0.4 luci1.0168-0171, luci2.0224-0228 completed
telluric Cresci C HIP102190 LUCI1+LUCI2 zj spec   luci1.0176-0177,luci2.0233-0234  
Antoniucci_11 Ves263 LUCI1+LUCI2 zj spec, Hkspec 0.8/0.8 luci1.0156-0163, luci2.0212-0219 completed
telluric Antoniucci_11 HD193984 LUCI1+LUCI2 zj spec, Hkspec   luci1.0190-0193, luci2.0247-0250 completed
Antoniucci_11 AGLL192 LUCI1+LUCI2 zj spec, Hkspec 0.6/0.6 luci1.0198-0201, luci2.0255-0258 completed
telluric Nisini Field2 HIP17971 LUCI1+LUCI2 J, H, K   luci1.0205-2017, luci2.0262-0273  
Nisini Field 2 LUCI1+LUCI2 J, H, K 1.75/3 h luci1.0226-0240, luci2.0282-0296 not every target are in the slits, maybe problem in the proper motion, or in field stop alignment, or both

Details

calibration, dark Cresci, telluric B,C
  • luci1.20180927.0141-0145, luci2.20180927.0118-0122 dit 5 ndit 3
  • luci1.20180927.0146-0150, luci2.20180927.0123-0127 dit 5 ndit 4

* Fiorentino*
  • luci2.0128-0132 N30 dit 10.0 ndit 30
  • luci2.0133-0137 N3.75 dit 2.74 ndit 10
  • luci2.0138-0142 N3.75 dit 10 ndit 10
  • luci2.0143-0145 N3.75 dit 10 ndit 20

02:30 problem with AO system

03:00 start Pinna program, target LO Peg luci2.0151-0154==check images. ==luci2.0155-0206 seeing variable between 1'' and 1.2'' *during the integration in K filter something happened in the dithering at K-band for this object, it wandered off the N30 field of view. the ISA fixed it, but there are like 5 exposures where the star is off to the side near the top.

** slew to Cresci program, seeing about 1.0''. send the telluric script, HIP92316
  • error in the preset
04:49 send again the script
  • error

05:06 send again lo script, now it works
  • acq luci1.0151-0152, luci2.0207-0208
  • 05:11 problem with the switch of mask in LUCI2
  • the mask in luci 2 had an error transferring from the grabber arm to the focal plane unit, Dave Thompson had to manually lock the mask into the focal plane.

05:49 problem fixed, the target now is too low and seeing is bed (variable between 1.1 and 1.9), we slew to the backup program Antoniucci, send the script for the telluric HIP116229 * 06:08: the preset failed, we send the script but the telescope hasn't received the information

06:11 send again the script
  • acq luci.0153-0155, luci2.0209-211
  • sci luci1.0156-0163, luci2.0212-0219 *seeing about 1.0'''

06:32 slew to the target V733
  • acq luci1.0164-0167, luci2.0220-0223
  • sci luci1.0168-0171, luci2.0224-0228 *06:45 error in LUCI *06:50 continue the integration, seeing 0.8''-0.9''

06:52 seeing seems stable (about 1.0''). we slew to Cresci program, telluric HIP102190
  • 6:55 check pointing
  • 07:04 load again the script
  • problem with the preset
  • 7:17 send the script
  • problem with the preset
  • 7:28 check pointing

07:30 send again the script for the telluric HIP102190
  • acq luci1.0173-0175, luci2.0229-0232
  • sci luci1.0176-0177,luci2.0233-0234

07:45 seeing around 2"-2.5", we can not continue with Cresci program
  • slew againt to Antoniucci program, target Ves263
  • acq luci1.0178-0181,luci2.0235-0238
  • sci luci1.0182-0186,luci2.0239-243
  • seeing 1.0"-1.1''

08:05 telluric HD193984
  • acq luci1.0187-0189, luci2.0244-0246
  • sci luci1.0190-0193, luci2.0247-0250
  • seeing 1.3"

08:22 target AGLL192
  • acq luci1.0194-0197, luci2.0251-0254
  • sci luci1.0198-0201, luci2.0255-0258
  • seeing 0.7"-0.8""
  • seeing 0.8'

08:54 slew to Cresci target C
  • error on pointing check
  • 09:07 we can collimated, but the seeing on the guide star is about 2.3" (problem in that particular part of the sky?with the guide star?), we change target

09:08 we slew to Nisini program because the targets of the other programs at higher ranking are too near to the moon (less the 30 degree)
  • 9:10 check pointing

09:17 load the script of the telluric for Nisini HIP17971
  • seeing 1.4-1.6"
  • acq luci1.0202-0204, luci2.0259-0261=
  • sci luci1.0205-2017, luci2.0262-0273
  • Seeing variable between 1" and 1.6"
  • 09:43 error in the grating guider
  • 09:55 continue the integrations, seeing about 1.0'''

10:01 field2 Nisini
  • acq luci1.0218-2020, luci2.0274-0276
  • luci1 stuck during the slit centering during the acquisition, we take by hand the image (slit+field), but somethings was wrong, we start again the acquisition
  • acq luci1.0222-0225, luci2.0277-0281
  • sci luci1.0226-0240, luci2.0282-0296

-- %USERSIG{RobertaCarini - 2018-09-26}%

Comments

 
Topic revision: r8 - 08 Nov 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback